Circular No. 8486

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005ah L. A. G. Monard, Pretoria, S. Africa, reports his discovery of an appar-

ent supernova (mag ∼ 17.3 ± 0.2) on unfiltered ST-7 CCD images taken on Feb. 10.775 UT (confirmed on Feb. 12.758 at mag ∼ 17.4 ± 0.2) with a 0.30m Schmidt-Cassegrain reflector. SN 2005ah is located at α = 6^{h}00^{m}45^{s}.77, δ = −58^{o}35′20′′.8 (equinox 2000.0), which is 27′′ west and 8′′ north of the nucleus of a galaxy in the cluster Abell S560. Nothing is visible at this position on an image from the Digitized Sky Survey (limiting red mag 20.5) or on an image taken by Monard on Jan. 4.883 (limiting red mag 18.8).

SUPERNOVA 2005ai IN NGC 2314 Further to IAUC 8470, T. Puckett and T. Orff report the discovery

of an apparent supernova (mag 15.9) on an unfiltered CCD frame taken with the 0.60-m automated supernova patrol telescope on Feb. 12.23 UT (and confirmed on frames taken by D. George, Ottawa, ON, with a 0.35-m reflector on Feb. 13.13 at mag 15.9; and by T. Crowley, Chiefland, FL, with a 0.30-m reflector on Feb 13.14, also at mag 15.9). The new object is located at α = 7^{h}10^{m}32^{s}.76, δ = +75^{o}21′29′′.1 (equinox 2000.0), which is 1′′.0 east and 113′′ north of the center of NGC 2314. Nothing was present at this position on images taken by Puckett on 2002 Apr. 1 and 2003 Apr. 27 (limiting mag ∼ 20.0).

SUPERNOVAE 2004gw, 2005T, AND 2005ae A. V. Filippenko and R. J. Foley, University of California, Berkeley,

report that inspection of CCD spectra (range 300–920 nm), obtained on Feb. 12 UT with the Keck I 10-m telescope (+ LRIS), shows that SN 2005ae (IAUC 8480) is probably of type IIb. The prominent He I series is indicative of a type-Ib supernova; however, there is an additional absorption line that corresponds to Hα at a velocity similar to that of the He I lines, which indicates that the progenitor had a low-mass hydrogen envelope. SN 2004gw is indeed of type Ia, as speculated by Foley et al. (IAUC 8465), and not of type Ic. Extremely noisy spectra of SN 2005T (IAUC 8473) reveal it to be of type II, with broad Hα and Ca II emission.

SUPERNOVA 2005W IN NGC 691 Unfiltered CCD magnitudes by E. Prosperi, Larciano, Italy: Feb. 3.770

UT, 14.8; 6.808, 14.4; 10.759, 14.3.

2005 February 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green